If I were to drive 160 miles traveling at how long would it take if I were to drive at 65mph? 75mph? and 85mph?
Also what formula would you use to solve this problem

Answer by Alan3354(69443)

If I were to drive 160 miles traveling at how long would it take if I were to drive at 65mph? 75mph? and 85mph?
-----------
d = rt
t = d/r = 160/r
------------
t = 160/65
t =~ 2.46 hours
-------------
You can do the others